OpenStudy (anonymous):

Simplify the radical. √w^6/16

OpenStudy (campbell_st):

is the question \[\sqrt{\frac{w^6}{16}}\]

OpenStudy (campbell_st):

or is it \[\frac{\sqrt {w^6}}{16}\]